Mixture of MgCO3​ & NaHCO3​ on strong heating gives CO2​ & H2​O in 3:1 mole ratio. The weight % of NaHCO3​ present in the mixture is :

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student

a. In the given reaction

MgCO3 = MgO + CO3

2NaHCO3= Na2O + H2O + 2CO2

b.In mixture of NaHCO3,MgCO3

Given that the mole ratio of CO2 & H2O is 3:1

c.That mean by given equation the mole ratio of NaHCO3 ,MgCO3 is 1:2

d. The molecular weight of NaHCO3 is 84 g/mole

e.The molecular weight of MgCO3 is 84.31 g/mole

f.Weight of 2 mole NaHCO3= 168 g

g. Let ,weight of mixture is= 168+84.3

= 252.3

h.Weight % of NaHCO3 in mixture

= (168/252.3)*100%

= 66.587%

thanks and regards